OpenStudy (anonymous):

A man invests $20,000 to yield $1,280 interest annually.If part of his investment pays 6% annually and the rest of his investment pays 7% annually,how much does he invest at each rate?

OpenStudy (anonymous):

we need a variable for the amount invested at say 7% you pick it

OpenStudy (anonymous):

it doesn't matter what you pick, so pick any letter you like

OpenStudy (anonymous):

idk that's the problem i don't understand the question

OpenStudy (anonymous):

i will walk you through it

OpenStudy (anonymous):

we need two numbers, the amount invested at 7% and the amount invested at 6% to do a problem like this, you need to have a variable representing some number you want, then you need an equation to solve but the first step is to make a variable you pick it

OpenStudy (anonymous):

pick some letter to represent the amount invested at 7% interest

OpenStudy (anonymous):

you can pick such your smart and you know what your doing

OpenStudy (anonymous):

there are 26 letters in the alphabet, you do not need to be smart to pick one of them

OpenStudy (anonymous):

just asking are you a boy or girl?

OpenStudy (anonymous):

also there is the greek alphabet, you can pick a letter from there too the indians used colors for there variables

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

why don't you pick \(x\) that is what the math teachers usually use

OpenStudy (anonymous):

ok \(x\) it is then if the amount invested at 7% is \(x\) and the total is \(20,000\) then the amount invested at 6% is \(20,000-x\)

OpenStudy (anonymous):

then

OpenStudy (misty1212):

7% of \(x\) is \(0.07x\) right? at 6% of \(20,,000-x\) is \(.06(20,000-x)\)

OpenStudy (anonymous):

yes, and since the total interest is 1280 you know that \[.07x+.06(20,000-x)=1280\]

OpenStudy (anonymous):

final job is to solve that for \(x\)

OpenStudy (misty1212):

yeah but decimals are so annoying

OpenStudy (anonymous):

k lets get rid of them by multiplying both sides by 100 to clear the decimals

OpenStudy (misty1212):

oh ok, then we can just solve \[7x+6(20,000-x)=128,000\]

OpenStudy (anonymous):

yeah if you multiply both sides by 100 the decimal goes, and of course you have to multiply \(1280\times 100\) to get \(128,000\)

OpenStudy (misty1212):

now it should be easy right? distribute the 6 then combine like terms then you can solve for \(x\) in one step

OpenStudy (anonymous):

so x is suppose to equal to?

OpenStudy (misty1212):

you still have to find \(x\) by solving \[7x+6(20,000-x)=128,000\]

OpenStudy (anonymous):

i mean how much does he invest at each rate?

OpenStudy (misty1212):

solve for \(x\) and you will know how much is invested at 7% do you know how to do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its all too confusing

OpenStudy (misty1212):

then skip it you need to know how to solve basic linear equations first before you can worry about complicated word problems

OpenStudy (anonymous):

but its homework and i need to finish it,its due tomorrow that's why i need help to find what x is

terenzreignz (terenzreignz):

Well, let's see what we can do ^^ Do you know what "distribute" means in this context? \[7x+\color{red}6(\color{blue}{20,000}-\color{green}x)=128,000\]

OpenStudy (misty1212):

it is too confusing go back and learn how to solve linear equations first then you need to know how to set up a word problem finally you can tackle something like this there is not royal road to algebra, you have to do it step by step

terenzreignz (terenzreignz):

Miss Tea scared her off D: Oh noes

OpenStudy (misty1212):

any one who starts of with "i don't get it" "this is too confusing" "this doesn't make sense" is not ready to begin a problem like this you could go ahead and give the answer, but it will make no sense at all look at all the work we did just to set up the equation need to go back and learn the basics, then we can tackle something like this

terenzreignz (terenzreignz):

Aww, there's no need to get so extreme ^^ Come on, @vickie11223 work with me here. What I want you to do is distribute the 6 to the two terms inside the parentheses. What that means is you multiply 6 to those two terms. Can you do that? \[7x+\boxed{6(20,000-x)}=128,000\] Focus on the stuff in the box first.

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):